OpenStudy (anonymous):

factor the following sum of two cubes: z^3+27. so far i have: (z+3)(x^2....

OpenStudy (skullpatrol):

That is a good start :-)

OpenStudy (luigi0210):

Do you know the formula?

OpenStudy (skullpatrol):

First write it as a "sum of two cubes."

OpenStudy (anonymous):

z^3+3^3

OpenStudy (anonymous):

is it (z+3)(x^2-3z+9)

OpenStudy (anonymous):

see a^3 +b^3 =(a+b)(a^2 -ab +b^2)

OpenStudy (anonymous):

yup correct

OpenStudy (anonymous):

do u want to factorise further...

OpenStudy (anonymous):

then if w be the complex cube root of unity then w^2+w=-1 and w^3=1 hence a^2 -ab +b^2=a^2 +(w^2+w)ab +b^2w^3 =a^2 +w^2ab+wab +b^2w^3 =a(a+bw^2) +bw(a+bw^2) =(a+bw^2)(a+bw)